OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. 2x2 + 7x + 6 = 0 double root real and rational root real and irrational root imaginary root

OpenStudy (dumbcow):

D = b^2 - 4ac D = 7^2 -4(2)(6) D = 1 2 rational roots